# Linear Bandits with Partially Observable Features

### v2

@@ Appendix D A Modified Algorithm for Time-Varying Observed Features
In this section, we define the problem of linear bandits with partially observable features under the setting where the observed features vary over time, describe our proposed method, and provide theoretical guarantees.
@@ D.1 Problem Formulation
Let $\mathbf{x}_{1,t},\ldots,\mathbf{x}_{K,t}$ denote the observed features and $\mathbf{u}_{1},\ldots,\mathbf{u}_{K}$ denote the unobserved features. Now the observed features arbitrary changes over $t$ but the unobserved features are fixed over time. When the algorithms selects an arm $a_{t}$, the reward is
@@
where the $\epsilon_{t}$ is the Sub-Gaussian noise that follows Assumption 1. The expected reward of each arm is sta over time, where MAB algorithms without using features are applicable to achieve $\widetilde{O}(\sqrt{KT})$ regret bound. When the observed features vary over time, the expected reward of each arm $\mathbb{E}[y_{a_{t}},t]=\langle\mathbf{x}_{a_{t},t},\boldsymbol{\theta}_{\star}^{(o)}\rangle+\langle\mathbf{u}_{a_{t}},\boldsymbol{\theta}_{\star}^{(u)}\rangle$ also arbitrarily changes over time and MAB algorithms suffer regret linear in $T$. To our knowledge, there is no other work that address this challenging setting.
@@ D.2 Proposed Method: Orthogonal Basis Augmentation
We address the problem by augmenting standard basis $\mathbf{e}_{1},\ldots,\mathbf{e}_{K}$ in $\mathbb{R}^{K}$ to estimate bias caused by the unobserved features. Let $\widetilde{\mathbf{x}}_{a,t}:=\mathbf{e}_{a}^{\top}[\mathbf{X}_{t}\;\mathbf{e}_{1}\;\cdots\;\mathbf{e}_{K}]\in\mathbb{R}^{d+K}$ and let $\Delta_{a}:=\langle\mathbf{u}_{a},\theta_{\star}^{(u)}\rangle$ denote the bias stems from the latent features. Then,
$y_{a,t}$ $=\langle\mathbf{x}_{a,t}^{\top}\boldsymbol{\theta}_{\star}^{(o)}\rangle+\langle\mathbf{u}_{a,t}^{\top}\boldsymbol{\theta}_{\star}^{(u)}\rangle+\epsilon_{t}$
$=\langle\mathbf{e}_{a}^{\top}[\mathbf{X}_{t}\mathbf{e}_{1}\cdots\mathbf{e}_{K}],[\boldsymbol{\theta}_{\star}^{(o)}\Delta_{1}\cdots\Delta_{K}]\rangle+\epsilon_{t}.$
Therefore, applying the RoLF-Ridge algorithm to the new features $\widetilde{\mathbf{x}}_{a,t}:=\mathbf{e}_{a}^{\top}[\mathbf{X}_{t}\mathbf{e}_{1}\cdots\mathbf{e}_{K}]$ yields the following regret bound.
Theorem 6 (Regret bound for Ridge-RoLF-V with time varying observed features).
If observed features are vary over time, for $\delta\in(0,1)$, with probability at least $1-4\delta$, the cumulative regret of the proposed algorithm Ridge-RoLF-V using DR Ridge estimator is bounded by
@@
The proof is similar to that in Theorem 5 and we omit the proof. The rate of the regret bound is $\widetilde{O}(\sqrt{(d+K)T})$ and, to our knowledge, this is the first sublinear regret bound for the partially observable linear bandits (as well as misspecified linear bandits) with arbitral time-varying observed features.
@@ Appendix E Missing Proofs
E.1 Proof of Theorem 1
Throughout this paper, we consider a bandit problem where the agent observes only a subset of the reward-generating feature vector and cannot access or estimate the unobserved portion. If the agent uses online decision-making algorithms that rely solely on observed features, as defined in Definition 1, the resulting issue can be interpreted as a model misspecification. Therefore, in this theorem, we present a problem instance where “misspecified” algorithms, considering only observed features, may incur regret that grows linearly in $T$.
Following the statement of Theorem 1, we assume that $d=d_{u}=1$, which means $d_{z}=2$. Given the true feature set $\mathcal{Z}=\{[1,3]^{\top},[2,19/4]^{\top}\}$, let the first element of each vector is observed to the agent; while the second element remains unobserved. This results in $\mathbf{x}_{1}=x_{1}=1$, $\mathbf{x}_{2}=x_{2}=2$, $\mathbf{u}_{1}=u_{1}=3$ and $\mathbf{u}_{2}=u_{2}=19/4$. We set the true parameter as $\boldsymbol{\theta}_{\star}\in\mathbb{R}^{2}=[2,-1]^{\top}$, meaning $\boldsymbol{\theta}_{\star}^{(o)}=\theta_{\star}^{(o)}=2$ and $\boldsymbol{\theta}_{\star}^{(u)}=\theta_{\star}^{(u)}=-1$. Using the reward function from Section 3.2 and considering Assumption 1, the expected reward for each arm is given by
@@ E.1 Proof of Theorem 1
Plugging the values in, the true mean reward for each arm is directly computed as $\gamma_{1}=2-3=-1$ and $\gamma_{2}=4-19/4=-3/4$, which satisfies the assumption that its absolute value does not exceed 1 (Section 3.2), and since $\gamma_{1}<\gamma_{2}$, the arm 2 is the optimal action.
For brevity, we denote the latent reward components as $g_{1}:=u_{1}\theta_{\star}^{(u)}$ and $g_{2}:=u_{2}\theta_{\star}^{(u)}$, yielding that $g_{1}=-3$ and $g_{2}=-19/4$. Note that since $\gamma_{2}\neq 2\gamma_{1}$ and $|{g_{i}}|\geq 3>0$ for all $i\in\{1,2\}$, thus our problem setup satisfies the “large deviation” criterion in Definition 1 and Theorem 2 of Ghosh et al. (2017), by letting $l=3$ and $\beta=0$. Consequently, by applying the theorem, it follows that using OFUL in this problem instance results in linear regret with respect to $T$, i.e., $\Omega(T)$. Inspired by this theorem, we demonstrate that LinTS encounters the same issue.
@@
$\begin{split}\widehat{\boldsymbol{\theta}}_{t}&=(\mathbf{X}_{t}^{\top}\mathbf{X}_{t}+\lambda\mathbf{I}_{d})^{-1}(\mathbf{X}_{t}^{\top}\mathbf{Y}_{t})\\ &=(\mathbf{X}_{t}^{\top}\mathbf{X}_{t}+\lambda\mathbf{I}_{d})^{-1}(\mathbf{X}_{t}^{\top}(\mathbf{X}_{t}\boldsymbol{\theta}_{\star}^{(o)}+\mathbf{g}_{t}+\boldsymbol{\epsilon}_{t}))\\ &=\boldsymbol{\theta}_{\star}^{(o)}-\lambda\mathbf{V}_{t}^{-1}\boldsymbol{\theta}_{\star}^{(o)}+\mathbf{V}_{t}^{-1}\mathbf{X}_{t}^{\top}\mathbf{g}_{t}+\mathbf{V}_{t}^{-1}\mathbf{X}_{t}^{\top}\boldsymbol{\epsilon}_{t},\end{split}$ (14)
@@
For this problem instance, since $d=1$Eq. 14 is equivalent to:
@@
where we assume $\lambda=1$. Note also that we denote $\widehat{\boldsymbol{\theta}}_{t}$ and $\boldsymbol{\theta}_{\star}^{(o)}$ by $\widehat{\theta}_{t}$ and $\theta_{\star}^{(o)}$, respectively, since both are scalars. Hence, the estimation error is computed as:
@@
where $\underline{g}=\min\{g_{1},g_{2}\}=-19/4$, which implies that $\sum_{\tau=1}^{t}x_{a_{\tau}}g_{a_{\tau}}=\Theta(t)$. For the denominator, $\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1$, as it grows at a rate of $O(t)$, implying that the second term of the right-hand side in Eq. 15 is $\Theta(1)$, and that $\widehat{\theta}_{t}$ is not consistent since it does not converge to $\theta_{\star}^{(o)}$ as $t\to\infty$.
For arm 2, which is optimal, to be selected in round $t+1$ under LinTS, the condition $x_{2}\widetilde{\theta}_{t}\geq x_{1}\widetilde{\theta}_{t}$ must hold, where $\widetilde{\theta}_{t}\sim\mathcal{N}\left(\widehat{\theta}_{t},\frac{v^{2}}{\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1}\right)$. Given the assumptions that $x_{1}=1$ and $x_{2}=2$, arm 2 is selected whenever $\widetilde{\theta}_{t}\geq 0$. Thus, for arm 1 to be chosen, we require $\widetilde{\theta}_{t}<0$. We will show that the probability of $\widetilde{\theta}_{t}<0$ does not diminish sufficiently to be ignored even the agent plays sufficiently large amount of time. To clarify, let us define two events $E_{\widetilde{\theta}}:=\{\widetilde{\theta}_{t}\geq 0\}$ and $E_{\hat{\theta}}:=\{\widehat{\theta}_{t}\geq 0\}$. We revisit Eq. 14 as follows:
@@
$\leq\theta_{\star}^{(o)}+\frac{g_{1}N_{1}+2g_{2}N_{2}}{N_{1}+4N_{2}+1}+\frac{\max_{a\in\{1,2\}}x_{a}}{N_{1}+4N_{2}+1}\sum_{\tau=1}^{t}\epsilon_{\tau}$
$=\theta_{\star}^{(o)}+\frac{g_{1}N_{1}+2g_{2}N_{2}}{N_{1}+4N_{2}+1}+\frac{2}{N_{1}+4N_{2}+1}\sum_{\tau=1}^{t}\epsilon_{\tau},$ (16)
where for the second term of Eq. 16, for $t\geq 19$,
@@
which is upper bounded by $-{9}/{4}$. This bound is followed by:
$\begin{split}\widehat{\theta}_{t}&\leq\theta_{\star}^{(o)}+\frac{g_{1}N_{1}+2g_{2}N_{2}}{N_{1}+4N_{2}+1}+\frac{2}{N_{1}+4N_{2}+1}\sum_{\tau=1}^{t}\epsilon_{\tau}\\ &\leq 2-\frac{9}{4}+\frac{2}{t+3N_{2}+1}\sum_{\tau=1}^{t}\epsilon_{\tau}\\ &=-\frac{1}{4}+\frac{2}{t+3N_{2}+1}\sum_{\tau=1}^{t}\epsilon_{\tau}\\ &\leq-\frac{1}{4}+\frac{2}{t}\sum_{\tau=1}^{t}\epsilon_{\tau}.\end{split}$ (17)
@@
Since $\epsilon_{\tau}$ is an IID sub-Gaussian random variable for all $\tau\in[t]$, by applying Hoeffding inequality we obtain:
$\mathbb{P}(\widehat{\theta}_{t}>0)\leq\exp\left(-\frac{t}{128\sigma^{2}}\right).$
Given this, we now bound the probability of the event $E_{\widetilde{\theta}}$:
$\mathbb{P}(E_{\widetilde{\theta}})$ $=\mathbb{P}(E_{\widetilde{\theta}}\cap E_{\hat{\theta}})+\mathbb{P}(E_{\widetilde{\theta}}\cap E_{\hat{\theta}}^{c})$
@@
Since the second term of Eq. 18 is calculated under a Gaussian distribution, which does not exceed $1/2$ for all $t\in[T]$,
$\mathbb{P}(E_{\widetilde{\theta}}^{c})\geq\frac{1}{2}-\exp\left(-\frac{t}{128\sigma^{2}}\right).$
Note that the total decision horizon $T>256\sigma^{2}\log(1/\delta^{\prime})$. For any $t>128\sigma^{2}\log(1/\delta^{\prime})$, we have $\mathbb{P}(E_{\widetilde{\theta}}^{c})\geq 1/2-\delta^{\prime}$. This implies that for the rounds more than $T/2$, the suboptimal arm is expected to be played at least $(1/2-\delta^{\prime})T/2$ times for any $\delta^{\prime}\in(0,1/2)$, thus incurring
@@
For OFUL, we also present another analysis that requires no assumption such that the suboptimal arm is played for initial $t$ rounds, which is taken in Theorem 2 of Ghosh et al. (2017). The optimal arm, arm 2, is selected when
@@
where $\widehat{\theta}_{t}$ is the same ridge estimator as in LinTS. The inequality Eq. 19 is equivalent to $\widehat{\theta}_{t}>(1+\sum_{\tau=1}^{t-1}x_{a_{\tau}}^{2})^{-1/2}$, which implies $\widehat{\theta}_{t}>1/\sqrt{2t}$. By Eq. 17,
@@
Thus, for $t\geq 128\sigma^{2}\log(2)$, the probability of selecting arm 2 is less than $1/2$ and for $T>256\sigma^{2}\log(2)$,

### v3

@@ Appendix D A Modified Algorithm for Time-Varying Observed Features
In this section, we define the problem of linear bandits with partially observable features under a setting where the observed features vary over time, describe our proposed method, and provide theoretical guarantees.
@@ D.1 Problem Formulation
Let $\mathbf{x}_{1,t},\ldots,\mathbf{x}_{K,t}$ denote the observed features and $\mathbf{u}_{1},\ldots,\mathbf{u}_{K}$ denote the unobserved features. Now the observed features arbitrarily change over $t$ but the unobserved features are fixed over time. When the algorithm selects an arm $a_{t}$, the reward is
@@
where $\epsilon_{t}$ is Sub-Gaussian noise that follows Assumption˜1. The expected reward of each arm is stable over time, where MAB algorithms without using features are applicable for achieving a $\widetilde{O}(\sqrt{KT})$ regret bound. When the observed features vary over time, the expected reward of each arm $\mathbb{E}[y_{a_{t},t}]=\langle\mathbf{x}_{a_{t},t},\boldsymbol{\theta}_{\star}^{(o)}\rangle+\langle\mathbf{u}_{a_{t}},\boldsymbol{\theta}_{\star}^{(u)}\rangle$ also arbitrarily changes over time, and MAB algorithms suffer regret linear in $T$. To our knowledge, there is no other work that addresses this challenging setting.
@@ D.2 Proposed Method: Orthogonal Basis Augmentation
We address the problem by augmenting standard basis $\mathbf{e}_{1},\ldots,\mathbf{e}_{K}$ in $\mathbb{R}^{K}$ to estimate bias caused by the unobserved features. Let $\widetilde{\mathbf{x}}_{a,t}:=\mathbf{e}_{a}^{\top}[\mathbf{X}_{t}\;\mathbf{e}_{1}\;\cdots\;\mathbf{e}_{K}]\in\mathbb{R}^{d+K}$ and let $\Delta_{a}:=\langle\mathbf{u}_{a},\theta_{\star}^{(u)}\rangle$ denote the bias that stems from the latent features. Then,
$y_{a,t}$ $=\langle\mathbf{x}_{a,t}^{\top},\boldsymbol{\theta}_{\star}^{(o)}\rangle+\langle\mathbf{u}_{a,t}^{\top},\boldsymbol{\theta}_{\star}^{(u)}\rangle+\epsilon_{t}$
$=\langle\mathbf{e}_{a}^{\top}[\mathbf{X}_{t}\;\mathbf{e}_{1}\;\ldots\;\mathbf{e}_{K}],[\boldsymbol{\theta}_{\star}^{(o)}\;\Delta_{1}\;\ldots\;\Delta_{K}]\rangle+\epsilon_{t}.$
Therefore, applying the RoLF-Ridge algorithm to the augmented features $\widetilde{\mathbf{x}}_{a,t}:=\mathbf{e}_{a}^{\top}[\mathbf{X}_{t}\;\mathbf{e}_{1}\;\ldots\;\mathbf{e}_{K}]$ yields the following regret bound.
Theorem 6 (Regret bound for Ridge-RoLF-V with time-varying observed features).
If the observed features are vary over time, then for any $\delta\in(0,1)$, with probability at least $1-4\delta$, the cumulative regret of the proposed algorithm Ridge-RoLF-V using DR Ridge estimator is bounded by
@@
The proof follows similar arguments given in Theorem˜5 and is therefore omitted. The resulting regret bound achieves a rate of $\widetilde{O}(\sqrt{(d+K)T})$, which, to the best of our knowledge, is the first sublinear regret guarantee for partially observable linear bandits (as well as misspecified linear bandits) with arbitrarily time-varying observed features.
@@ Appendix E Missing Proofs
E.1 Proof of Theorem˜1
Throughout this paper, we consider a bandit problem where the agent observes only a subset of the reward-generating feature vector and cannot access or estimate the unobserved portion. If the agent uses online decision-making algorithms that rely solely on observed features, as defined in Definition˜1, the resulting issue can be interpreted as a model misspecification. Therefore, in this theorem, we present a problem instance where “misspecified” algorithms, considering only observed features, may incur regret that grows linearly in $T$.
Following the statement of Theorem˜1, we assume that $d=d_{u}=1$, which means $d_{z}=2$. Given the true feature set $\mathcal{Z}=\{[1,3]^{\top},[2,19/4]^{\top}\}$, let the first element of each vector is observed to the agent; while the second element remains unobserved. This results in $\mathbf{x}_{1}=x_{1}=1$, $\mathbf{x}_{2}=x_{2}=2$, $\mathbf{u}_{1}=u_{1}=3$ and $\mathbf{u}_{2}=u_{2}=19/4$. We set the true parameter as $\boldsymbol{\theta}_{\star}\in\mathbb{R}^{2}=[2,-1]^{\top}$, meaning $\boldsymbol{\theta}_{\star}^{(o)}=\theta_{\star}^{(o)}=2$ and $\boldsymbol{\theta}_{\star}^{(u)}=\theta_{\star}^{(u)}=-1$. Using the reward function from Section˜3.2 and considering Assumption˜1, the expected reward for each arm is given by
@@ E.1 Proof of Theorem˜1
Plugging the values in, the true mean reward for each arm is directly computed as $\gamma_{1}=2-3=-1$ and $\gamma_{2}=4-19/4=-3/4$, which satisfies the assumption that its absolute value does not exceed 1 (Section˜3.2), and since $\gamma_{1}<\gamma_{2}$, arm 2 is the optimal action. We further assume that the total learning horizon $T>256\sigma^{2}\max\{\log 2,\log(1/\delta^{\prime})\}$ for any $\delta^{\prime}\in(0,1/2)$, where $\sigma$ is the sub-Gaussian parameter of the reward noise.
For brevity, we denote the latent reward components as $g_{1}:=u_{1}\theta_{\star}^{(u)}$ and $g_{2}:=u_{2}\theta_{\star}^{(u)}$, yielding $g_{1}=-3$ and $g_{2}=-19/4$. Since $\gamma_{2}\neq 2\gamma_{1}$ and $\lvert g_{i}\rvert\geq 3>0$ for all $i\in\{1,2\}$, our problem setup satisfies the “large deviation” criterion in Definition 1 and Theorem 2 of Ghosh et al. (2017), with $l=3$ and $\beta=0$. Applying the theorem, it follows that OFUL (Abbasi-yadkori et al., 2011) suffers linear regret in this problem instance, i.e., $\Omega(T)$. Motivated by this result, we further show that LinTS (Agrawal & Goyal, 2013) is similarly affected.
@@
$\widehat{\boldsymbol{\theta}}_{t}$ $=(\mathbf{X}_{t}^{\top}\mathbf{X}_{t}+\lambda\mathbf{I}_{d})^{-1}(\mathbf{X}_{t}^{\top}\mathbf{Y}_{t})$
$=(\mathbf{X}_{t}^{\top}\mathbf{X}_{t}+\lambda\mathbf{I}_{d})^{-1}(\mathbf{X}_{t}^{\top}(\mathbf{X}_{t}\boldsymbol{\theta}_{\star}^{(o)}+\mathbf{g}_{t}+\boldsymbol{\epsilon}_{t}))$
$=\boldsymbol{\theta}_{\star}^{(o)}-\lambda\mathbf{V}_{t}^{-1}\boldsymbol{\theta}_{\star}^{(o)}+\mathbf{V}_{t}^{-1}\mathbf{X}_{t}^{\top}\mathbf{g}_{t}+\mathbf{V}_{t}^{-1}\mathbf{X}_{t}^{\top}\boldsymbol{\epsilon}_{t},$ (14)
@@
For this problem instance, since d=1d=1Eq.˜14 is equivalent to:
@@
where we assume $\lambda=1$. Note that we also denote $\widehat{\boldsymbol{\theta}}_{t}$ and $\boldsymbol{\theta}_{\star}^{(o)}$ by $\widehat{\theta}_{t}$ and $\theta_{\star}^{(o)}$, respectively, since both are scalars. Hence, the estimation error is computed as:
@@
where $\underline{g}=\min\{g_{1},g_{2}\}=-19/4$, which implies that $\sum_{\tau=1}^{t}x_{a_{\tau}}g_{a_{\tau}}=\Theta(t)$. For the denominator, $\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1$, which grows at a rate of $O(t)$, implying that the second term of the right-hand side in Eq.˜15 is $\Theta(1)$, and that $\widehat{\theta}_{t}$ is not consistent since it does not converge to $\theta_{\star}^{(o)}$ as $t\to\infty$.
For arm 2, which is optimal, to be selected in round $t+1$ under LinTS, the condition $x_{2}\widetilde{\theta}_{t}\geq x_{1}\widetilde{\theta}_{t}$ must hold, where $\widetilde{\theta}_{t}\sim\mathcal{N}\left(\widehat{\theta}_{t},v^{2}(\sum_{\tau=1}^{t}x_{a_{\tau}}^{2}+1)^{-1}\right)$. Given the assumptions that $x_{1}=1$ and $x_{2}=2$, arm 2 is selected whenever $\widetilde{\theta}_{t}\geq 0$. Thus, for arm 1 to be chosen, we require $\widetilde{\theta}_{t}<0$. We will show that the probability of $\widetilde{\theta}_{t}<0$ does not diminish sufficiently to be ignored, even when the agent plays for a sufficiently large amount of time. To clarify, let us define the two events $E_{\widetilde{\theta}}:=\{\widetilde{\theta}_{t}\geq 0\}$ and $E_{\hat{\theta}}:=\{\widehat{\theta}_{t}\geq 0\}$. We revisit Eq.˜14 as follows:
@@
$\leq\theta_{\star}^{(o)}+\frac{g_{1}N_{1}+2g_{2}N_{2}}{N_{1}+4N_{2}+1}+\frac{2}{N_{1}+4N_{2}+1}\sum_{\tau=1}^{t}\epsilon_{\tau},$ (16)
where the last inequality holds since $\sum_{\tau=1}^{t}x_{a_{\tau}}\epsilon_{\tau}\leq\max_{a\in\{1,2\}}x_{a}\sum_{\tau=1}^{t}\epsilon_{\tau}$ and $\max_{a\in\{1,2\}}x_{a}=2$. For the second term of Eq.˜16, for $t\geq 19$,
@@
which is upper bounded by 9/4-{9}/{4}. This bound is followed by:
$\begin{split}\widehat{\theta}_{t}&\leq\theta_{\star}^{(o)}+\frac{g_{1}N_{1}+2g_{2}N_{2}}{N_{1}+4N_{2}+1}+\frac{2}{N_{1}+4N_{2}+1}\sum_{\tau=1}^{t}\epsilon_{\tau}\\ &\leq 2-\frac{9}{4}+\frac{2}{t+3N_{2}+1}\sum_{\tau=1}^{t}\epsilon_{\tau}\\ &\leq-\frac{1}{4}+\frac{2}{t}\sum_{\tau=1}^{t}\epsilon_{\tau}.\end{split}$ (17)
@@
Since $\epsilon_{\tau}$ is an IID sub-Gaussian random variable for all $\tau\in[t]$, by applying Hoeffding inequality we obtain: $\mathbb{P}(\widehat{\theta}_{t}>0)\leq\exp\left(-t/128\sigma^{2}\right)$. Given this, we now bound the probability of the event $E_{\widetilde{\theta}}$:
[…]
@@
Since the second term of Eq.˜18 is calculated under a Gaussian distribution, thus its value does not exceed $1/2$ for all $t\in[T]$, it follows that $\mathbb{P}(E_{\widetilde{\theta}}^{c})\geq 1/2-\exp\left(-{t}/128\sigma^{2}\right).$ Note that the total decision horizon $T>256\sigma^{2}\log(1/\delta^{\prime})$, thus for any $t>128\sigma^{2}\log(1/\delta^{\prime})$, we have $\mathbb{P}(E_{\widetilde{\theta}}^{c})\geq 1/2-\delta^{\prime}$. This implies that for rounds beyond $T/2$, the suboptimal arm will be played at least $(1/2-\delta^{\prime})T/2$ times for any $\delta^{\prime}\in(0,1/2)$, thus incurring
@@
For OFUL, we also present another analysis that does not require the assumption that the suboptimal arm is played for initial $t$ rounds, which is taken in Theorem 2 of Ghosh et al. (2017). The optimal arm, arm 2, is selected when
@@
where $\widehat{\theta}_{t}$ is the same ridge estimator as in LinTS. The inequality Eq.˜19 is equivalent to $\widehat{\theta}_{t}>(1+\sum_{\tau=1}^{t-1}x_{a_{\tau}}^{2})^{-1/2}$, which implies $\widehat{\theta}_{t}>1/\sqrt{2t}$. By Eq.˜17,
@@
Thus, for $t\geq 128\sigma^{2}\log 2$, the probability of selecting arm 2 is less than $1/2$ and for $T>256\sigma^{2}\log 2$,